OpenStudy (anonymous):

a barrel contains 183 gallon of gas, and it is being drained at a constant rate of 9 gallons a minute. Write equation that models the numberes of gallon, g, after t minutes. The equation that model the given situation is ??

Parth (parthkohli):

The original amount is 183, and it is decreasing at 9 per minute.

OpenStudy (anonymous):

correct I thought that you would divide 183/9 to get the situation

Parth (parthkohli):

In 1 minute, the barrel would have decreased by \(9 \times 1\) gallons. In 2, would've decreased by \(9 \times 2\) gallons.

Parth (parthkohli):

In 1 minute, the barrel would have contained \(\rm 183 - (9 \times 1)\) gallons. In 2 minutes, the barrel would have contained \(\rm 183 - (9 \times 2)\) gallons.

OpenStudy (anonymous):

so the first one would be 183-9=174 and the second one is 183-18=165 do I keep going

Parth (parthkohli):

But you have to model an equation. What would it be?

OpenStudy (anonymous):

183-(9*1)

OpenStudy (anonymous):

right

OpenStudy (anonymous):

can someone please help me do not understand how to do equation

Parth (parthkohli):

Hmm, you have to model an equation which represents the number of gallons left in 't' minutes.

OpenStudy (anonymous):

I am sorry I do not understand so you are telling me if I have 183 gallons and 9 are being taking away every minute so would it be 183-(9*x)